OpenStudy (anonymous):

Solve –3(2x – 1) = x + 10. You must show all work.

OpenStudy (anonymous):

–3(2x – 1) = x + 10 -6x + 3 = x + 10 -6x - x = 10 - 3 -7x = 7 Can you finish it?